OpenStudy (anonymous):

Solve for x and y. 3x - 4y = -10 5x + 8y = -2

OpenStudy (anonymous):

5(3x-4y)=-50 3(5x+8y)=-6 15x-20y=-50 15x+24y=-6 subtract the equations and solve for y then plug the value of y in one of the equations to find x

OpenStudy (saifoo.khan):

Try them sis.